# DualMap: Online Open-Vocabulary Semantic Mapping for Natural Language Navigation in Dynamic Changing Scenes

### v3

@@ B-D Dynamic Setting
The added YCB dataset objects are also used in dynamic change query experiments. Following the two types of dynamic changes defined in Sec.V-A1, each object is moved three times for both in-anchor and cross-anchor changes. Fig. 7 uses scene 00829 as an example to illustrates how these changes are applied.
@@ C-A Feature Weights
TABLE IX: Ablation Study on Feature Weighting
@@
Regarding the weighting in Equation 1 in Sec. III-A2, we empirically selected the weights based on an ablation study conducted on both the Replica and ScanNet datasets. As shown in Table IX, assigning weights of 0.7 and 0.3 to $\mathbf{f}_{\text{image}}$ and $\mathbf{f}_{\text{text}}$, respectively, yields the best or near-best performance across all metrics.
@@ C-B Detailed Object Classification in Abstraction
TABLE X: Open-vocabulary 3D Semantic Segmentation and Efficiency
@@ D-A Results with Other CLIP Backbone
The results presented in Table X highlight the advantages of using the Mobile-CLIP backbone for semantic segmentation tasks. Notably, methods employing Mobile-CLIP demonstrate superior accuracy and faster performance compared to those utilizing ViT-H/14 backbone. Our proposed method not only exhibits robustness across various CLIP backbones but also consistently outperforms other methods. Additionally, substituting Mobile-CLIP for ViT-H/14 significantly reduces memory usage and time costs, making it more resource-efficient for semantic segmentation applications.
@@ E-A Runtime Decomposition
We break down the module-wise time cost on RTX 4090 Desktop in Table XI and Figure 12. The results show that over 57% of the time per frame is spent on model inference (highlighted in grey). This indicates that applying inference optimizations such as TensorRT acceleration or model quantization can significantly reduce runtime with minimal accuracy impact.
@@ E-B Experiments on Laptop and Different Resolutions
While our main experiments were conducted on an RTX 4090 GPU Desktop, we conducted additional experiments on a laptop equipped with an RTX 3080 Laptop GPU. The additional system evaluation on Replica dataset is shown in Table XII. The results show that, at the same input resolution (1200×680), the RTX 3080 Laptop achieves comparable accuracy with a  1.6× increase in latency (around 0.4s/frame), which remains acceptable for online applications. Furthermore, by reducing the input resolution to 640×360, the system achieves a good trade-off between accuracy and efficiency, with only a minor drop in accuracy (e.g., ~3% FmIoU loss) and comparable runtime performance to the results on the RTX 4090 Desktop. This suggests that appropriately adjusting input resolution offers a practical path for deployment on lower-cost devices.
TABLE XI: Time Decomposition on Replica with 4090
@@
TABLE XII: Accuracy and Efficiency Across GPU and Resolution Settings on Replica

### v4

@@ B-D Dynamic Setting
The added YCB dataset objects are also used in dynamic change query experiments. Following the two types of dynamic changes defined in Sec.V-A1, each object is moved three times for both in-anchor and cross-anchor changes, showing in Fig. 7. To facilitate reproducibility, detailed experiment results are presented in Tables IX and X.
TABLE IX: Detailed In-Anchor Experiment Results
00829 00848 00880
Type Success Type Success Type Success
0116.json
soup_can 1 scissors 1 soup_can 1
bowl 1 plate 1 bowl 1
plate 1 pitcher 0 pitcher 1
scissors 0 mug 1 plate 0
cracker_box 1 banana 1 cracker_box 0
pitcher 1 cracker_box 1 scissors 0
0117.json
soup_can 1 scissors 1 soup_can 1
bowl 1 plate 1 bowl 1
plate 1 pitcher 0 pitcher 1
scissors 0 mug 0 plate 0
cracker_box 1 banana 1 cracker_box 0
pitcher 0 cracker_box 1 scissors 1
0118.json
soup_can 0 scissors 0 soup_can 1
bowl 1 plate 1 bowl 1
plate 1 pitcher 0 pitcher 1
scissors 0 mug 0 plate 0
cracker_box 1 banana 1 cracker_box 0
pitcher 0 cracker_box 1 scissors 1
Summary
Total 12/18 Total 12/18 Total 11/18
TABLE X: Detailed Cross-Anchor Experiment Results
00829 00848 00880
Type Success Type Success Type Success
0129-1.json
bowl 1 cracker_box 1 soup_can 1
cracker_box 0 scissors 1 pitcher 0
soup_can 1 mug 0 bowl 1
plate 1 pitcher 1 plate 0
pitcher 1 plate 1 scissors 1
scissors 0 banana 0 cracker_box 1
0128-2.json
bowl 1 cracker_box 1 soup_can 1
cracker_box 0 scissors 1 pitcher 0
soup_can 0 mug 0 bowl 1
plate 0 pitcher 0 plate 1
pitcher 1 plate 1 scissors 0
scissors 1 banana 1
0128-1.json
bowl 1 cracker_box 0 soup_can 0
cracker_box 0 scissors 1 pitcher 1
soup_can 1 mug 1 bowl 1
plate 1 pitcher 0 plate 1
pitcher 0 scissors 1 scissors 0
scissors 0 banana 1 cracker_box 0
Summary
Total 10/18 Total 12/18 Total 10/17
@@ C-A Feature Weights
TABLE XI: Ablation Study on Feature Weighting
@@
Regarding the weighting in Equation 1 in Sec. III-A2, we empirically selected the weights based on an ablation study conducted on both the Replica and ScanNet datasets. As shown in Table XI, assigning weights of 0.7 and 0.3 to $\mathbf{f}_{\text{image}}$ and $\mathbf{f}_{\text{text}}$, respectively, yields the best or near-best performance across all metrics.
@@ C-B Detailed Object Classification in Abstraction
TABLE XII: Open-vocabulary 3D Semantic Segmentation and Efficiency
@@ D-A Results with Other CLIP Backbone
The results presented in Table XII highlight the advantages of using the Mobile-CLIP backbone for semantic segmentation tasks. Notably, methods employing Mobile-CLIP demonstrate superior accuracy and faster performance compared to those utilizing ViT-H/14 backbone. Our proposed method not only exhibits robustness across various CLIP backbones but also consistently outperforms other methods. Additionally, substituting Mobile-CLIP for ViT-H/14 significantly reduces memory usage and time costs, making it more resource-efficient for semantic segmentation applications.
@@ E-A Runtime Decomposition
We break down the module-wise time cost on RTX 4090 Desktop in Table XIII and Figure 12. The results show that over 57% of the time per frame is spent on model inference (highlighted in grey). This indicates that applying inference optimizations such as TensorRT acceleration or model quantization can significantly reduce runtime with minimal accuracy impact.
@@ E-B Experiments on Laptop and Different Resolutions
While our main experiments were conducted on an RTX 4090 GPU Desktop, we conducted additional experiments on a laptop equipped with an RTX 3080 Laptop GPU. The additional system evaluation on Replica dataset is shown in Table XIV. The results show that, at the same input resolution (1200×680), the RTX 3080 Laptop achieves comparable accuracy with a  1.6× increase in latency (around 0.4s/frame), which remains acceptable for online applications. Furthermore, by reducing the input resolution to 640×360, the system achieves a good trade-off between accuracy and efficiency, with only a minor drop in accuracy (e.g., ~3% FmIoU loss) and comparable runtime performance to the results on the RTX 4090 Desktop. This suggests that appropriately adjusting input resolution offers a practical path for deployment on lower-cost devices.
TABLE XIII: Time Decomposition on Replica with 4090
@@
TABLE XIV: Accuracy and Efficiency Across GPU and Resolution Settings on Replica